OpenStudy (anonymous):

Factor as much as possible x^2+15x+14

OpenStudy (anonymous):

x^2+210x

OpenStudy (anonymous):

You want to find two numbers a and b such that ab = 14 (the product is 14) and the sum a + b = 15. We can see a = 1 and b = 14 will do it for us. So, split 15x = 1x + 14x... x^2 + x + 14x + 14 = x(x + 1) = 14(x + 1) = (x + 14)(x + 1)